A patient is receiving 1328 mL of a 2 in 1 solution of 4 AA/D25W. Calculate the estimated osmolarity of this volume of solution.
Answer #1

To calculate osmolarity

a.First multiply grams of dextrose per litre by 5.

For 25%dextrose it is 250mg

250*5=1250ml

b.Multiply grams of protein per litre by 10

45=40g

40g*10=400mOsm

c.Add both a and b.

d Add 300 to 400 to the answer from c because vitamins and minerals contribute about 300 to 400mOsm/L

1650+300 to 400=1950 to 2050mOsm/L